OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (4, 0) and (2, 6)?

OpenStudy (anonymous):

y = -3x + 12

OpenStudy (anonymous):

wrong .

OpenStudy (anonymous):

oh i guess that's slope-intercept form

OpenStudy (anonymous):

these are the answers y = 4x - 2 y = 2x - 4 y = -3(x - 4) y = 3(x + 4)

OpenStudy (anonymous):

haha well -3(x-4) is the same as what i wrote